Mathematics majors generally have one of the highest starting salaries. Suppose the average starting salary for mathematics majors is $48000 with a standard deviation of $6900. Find the probability that 35 mathematics majors have an average starting salary less than $48079 thousand.
Answer #1
for normal distribution z score =(X-μ)/σx
here mean=       μ= 48000
std deviation   =σ= 6900
sample size       =n= 35
std error=σ=σ/√n= 1166.31287
probability =P(X<48079)=(Z<(48079-48000)/1166.313)=P(Z<0.07)=0.5279